# HallE-Switch: Rethinking and Controlling Object Existence Hallucinations in Large Vision Language Models for Detailed Caption

### v1

@@ 3 Hallucination Controlling
We introduce HallE-Switch, a LVLM designed to control the extent of parametric knowledge within detailed captions. For this purpose, we developed two datasets: the first captures solely contextual knowledge, while the second merges both contextual and parametric knowledge. Using these datasets, we integrated a control projector into the model to control parametric knowledge.
@@ 3.1 Data Generation
We present an approach to generating new data on MSCOCO (Sharma et al., 2018) source labels using GPT-4, follwing LLaVA’s methodology.
Grouping using RAM. We begin by passing MSCOCO’s ground truth objects utilizing the open vocabulary detector RAM (Zhang et al., 2023). This step aims to simulate the maximum visual granularity achievable by a vision encoder. RAM categorizes the data into sets: “grounded” (contextual group) and “not grounded” (parametric group).
Contextual Data Generation. For our first dataset, we adhere to LLaVA’s original detailed caption creation pipeline. This involves generating detailed captions using only objects from the contextual group. To do this, we feed MSCOCO source labels (including object classes, bounding boxes, and short captions) into GPT-4.
Parametric Joint Data Generation. The second dataset incorporates both contextual and parametric knowledge. Here, we begin with LLaVA’s original detailed captions and annotate objects from the parametric group with special tokens. Specifically, we enclose the ’undetected’ objects with brackets. Formally, if denotes the original image caption sentence and represents a set of non-grounded objects, our data processing can be represented as:
@@
3.2 Hallucination Switch
Inspired by LM-Switch (Han et al., 2023), we propose to address hallucination by learning a control parameter that serves as a “switching value”, e.g., for permitted imagination and for restricting imagination, as depicted in Figure 2(a). Let represent the LLM with fixed parameters. Its output word embedding is modified as , leading to the derived model as:
The learned can be regarded as the transformation from a generic word space to the object sensitive word space, where word-word semantic correspondence is optimized to object correspondence, and governs the intensity of imagining related objects.
@@ 3.2 Hallucination Switch
Inference. At the inference stage, can adopt any value within the interval . Specifically, an value of -1 corresponds to minimal reliance on parametric knowledge, whereas a value of 1 indicates a strong inclination towards such knowledge. Detailed theoretical explanation are elaborated upon in the Appendix.
@@ 4 Experiment
| Setting | LLM | Resolution | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object |
@@
| Switch | LLM | Resolution | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object |
@@ 4.1 Finetune on Parametric Joint Data
Before we present experiments on HallE-Switch, we show the upper-bound experiment result on how well the model can indicate parametric knowledge. We directly finetune the LLaVA model on parametric joint data. Intuitively, the model is trained on data indicating parametric knowledge. Its output should identify parametric knowledge accurately. Specifically, the model should put a bracket around every ”guessed” objects for indication of hallucination.
Therefore, we evaluate the object hallucination in three different settings: 1. Evaluation only on indicated objects: We do CCEval which evaluate CHAIR, CHAIR, and coverage on objects inside the bracket. The result should reflect a high level of hallucination. 2. Evaluation without indicated objects: We do CCEval which evaluate CHAIR, CHAIR, and coverage on objects not in the bracket. The result should reflect a low level of hallucination. 3. Evaluation with indicated objects. We do CCEval as following:
Evaluation only on indicated objects. The hallucination level for indicated objects, denoted as , is 63.90 for LLaVA and 62.31 for LLaVA. It is considerably higher than the baselines all other models. Concurrently, their coverage is 12.01 and 19.09 for LLaVA and LLaVA, respectively, which both of them significantly lower than the coverage of 33.58 for LLaVA baseline model. The experiment results indicate the object within special tokens has significantly higher hallucination rate and lower coverage rate which support our assumption that the object inside the indication tokens are objects indicated from parametric knowledge.
Evaluation without indicated objects. For objects outside of the special token scope, we found that hallucination is markedly reduced, CHAIR decreased from 82 to 57 which is 30.5% percent improvement and CHAIR decrease 32.4%, from 25.3 to 17.10, compared to the baseline. This suggests that the model is less prone to make erroneous assumptions for objects not marked by brackets.
Evaluation with indicated objects.: We observe a significant decline in the hallucination without any reduce in object coverage. LLaVA CHAIR improved from 25.30 to 14.00 which has 44.66% improvements. For LLaVA CHAIR improved from 16 to 9.86 which also has 38.38% improvements.
@@ 4.2 Hallucination Controlling
During model inference, we select 4 different , ranging from to . As shown in Table 8, we evaluate HallE-Switch and HallE-Switch model, which use LLaVA as backbone. For the 7B model, we train it by removing the special token in parametric joint data, showing that indication is not a necessary condition for switch to work. means the switch trained purely on contextual only data, which try to minimize the hallucination, where switch to maximize parametric knowledge output. The results show that as increase, CHAIR increase from 20.90 to 26.6, the coverage keeps at a similar level.
For the 13B model, we keep the indication inside the parametric joint data. The HallE-Switch achieves the best in object existence hallucination metric. With switch set to -1 and indication, we have versus baseline’s 64 and vs baseline’s 16, and the coverage of the model is not decreasing.
@@ 5 Related Work
Large Vision-Language Models (LVLMs).
The rapid advancements in Large Language Models (LLMs) (Touvron et al., 2023a; Chung et al., 2022; Touvron et al., 2023b; Anil et al., 2023; Driess et al., 2023; Scao et al., 2022) combined with a surge in open-source initiatives, such as GPTs, LLaMAs, MPT, and Vicuna, have paved the way for the emergence of extensive vision-language models Liu et al. (2023a); Goyal et al. (2017); Li et al. (2023b); Zhu et al. (2023); Sun et al. (2023); Ye et al. (2023); Bai et al. (2023); Chen et al. (2023); Peng et al. (2023) like GPT-4 OpenAI (2023). LLaVA introduced the concept of integrating a simple projector during LLM fine-tuning. Concurrently, BLIP2 (Li et al., 2023c)presented q-former-based LVLMs, with InstructBLIP (Dai et al., 2023) further refining the concept. Otter aims to improve OpenFlamingo’s Alayrac et al. ; Awadalla et al. (2023) directive adherence. MiniGPT-4 Zhu et al. (2023) highlights the potential of sophisticated LLMs and advocates for a unified layer to synchronize visual and text models, offering GPT4-like capabilities but with reduced computational demands. mPLUG-Owl Ye et al. (2023) suggests a two-step method: first aligning visual aspects, then refining the language model using techniques like LoRA. QWen-VL Bai et al. (2023) scales up LMM pre-training significantly and showcases stellar performance on various benchmarks.
@@
Before the emergence of LLMs, the NLP community characterized “hallucination” as the generation of content that either lacked coherence or deviated from its original source (Ji et al., 2023a). Hallucinations have been widely studied in traditional NLG (Ji et al., 2023b) tasks, including machine translation (Zhou et al., 2020; Lee et al., 2019), data-to-text (Rebuffel et al., 2021; Kasner & Dušek, 2022; Lee et al., 2022), summarization (Cao et al., 2022), dialogue (Dziri et al., 2022) and QA (Shuster et al., 2021). After LLMs significantly bring model performance to next stage, several works bring new meaning to hallucination as the generation of text that is erroneous, nonsensical, or detached from reality (Shi et al., 2023; Lin et al., 2021). For LVLMs, previous studies have been mainly focusing on object hallucination (Marino et al., 2019; MacLeod et al., 2017; Li et al., 2023a; d).
@@ 6 Conclusion
In summary, this study delves deep into the object hallucination phenomena within the detailed captions of LVLMs, advancing understanding of the accuracy and unwarranted inference in describing visual details. We introduce a novel and comprehensive evaluation method for object existence hallucination in detailed captions. We conduct an in-depth and component-wise analysis of LVLMs, meticulously examining each element that might result in hallucination. We further identify an alignment issue between the vision encoder and the instruction data. To alleviate such hallucination, we introduce controlling parameters over LVLMs to condition the inference of objects.
@@ Appendix A Appendix
A.1 Training Data Quality
@@
Image |
Caption |
GT Labels |
@@
The image depicts a cozy entertainment room featuring a flat screen TV situated against the wall. A dining table can be seen next to the TV, with some bottles and a cup placed on top of it. Multiple video game controllers, including Wii remotes, are also present on the dining table, suggesting that people are using the TV to play video games. To the far right of the scene, another remote is located close to the wall. A person appears to be sitting or standing behind the dining table, likely either watching the game or waiting for their turn to play. The assortment of bottles on the table suggests that the guests are enjoying drinks during their video game session. |
tv: 1 cup: 1 remote: 3 bottle: 4 person: 1 dining table: 1 |
@@
In the image, an orange mass transit trolley is making its way through a city. A person is crossing the street in front of the trolley while holding a garbage bag, appearing to be cautious about the approaching vehicle. Another person is standing close to the person crossing the street, and there are two more individuals nearby. In the scene, various vehicles surround the trolley, including cars, buses, and trucks. One of the cars is parked right behind the trolley, while another is situated farther back. Two buses can be seen, with one staying behind the trolley and the other on the right side of it. A truck is also present at the far left side of the scene. A woman nearby is holding a handbag, completing the busy urban setting. |
car: 3 bus: 2 truck: 1 train: 1 person: 5 handbag: 1 |
@@
The image shows a table in a restaurant, which appears to have been recently used for a meal. The table is set with four place settings, including white dishes. On the table, there are multiple cups, a bowl, and a mix of silverware like forks and spoons laid out. Some of the dishes look dirty, with used napkins and eating utensils scattered around. In the background, two people can be seen, one situated on the left side and another on the right side behind the table. There are also chairs located near the table, with one chair positioned close to the left side and another chair closer to the right side of the frame. |
cup: 5 spoon: 3 fork: 2 person: 2 chair: 2 bowl: 1 dining table: 1 |
A.2 Prompt
@@ A.4 Experiment Settings
Hallucination switch only involves only finetuning a linear layer added before layer. We freeze all other layers and only finetune switch layer with 3 epochs. The dataset is generated data with associated value. We have 10K contextual only detailed caption data and 23K parametric joint detailed caption data.
A.5 Theoretical Explanations
HallE-Switch Formulation
HallE-Switch follow LLaVA’s training strategy, which freezes vision encoder and language model, only finetune the projector for the first stage. The visual encoder transfer image to visual features:

### v3

@@ 3 Hallucination Controlling
Building on our prior finding that parametric knowledge can lead to hallucination, controlling or signaling the model’s use of this knowledge enables us to manage its tendency for hallucination. By incorporating a controller layer, we directly influence the output distribution linked to parametric knowledge. To this end, we developed HallE-Control, a LMM designed to control the extent of parametric knowledge within detailed captions. For this purpose, we developed two datasets to train the controller: the first captures solely contextual knowledge, while the second merges both contextual and parametric knowledge. Additionally, we can make the model signals the parametric knowledge by marking unrecognised objects in the finetuning dataset.
@@ 3.1 Data Generation
Grouping using RAM. We begin by passing MSCOCO’s ground truth objects to the open vocabulary detector RAM (Zhang et al., 2023b). RAM categorizes the objects into two groups: “grounded” (detected objects as contextual group) and “omitted” (not detected objects as parametric group). This step aims to simulate the maximum visual granularity achievable by a vision encoder in LMMs.
Contextual Data Generation. Our first dataset involves generating detailed captions using only objects from the contextual group. To do this, we feed MSCOCO source labels (including object classes, bounding boxes, and short captions) into GPT-4. We adhere to LLaVA’s caption creation pipeline and provide prompt in Appendix.
Parametric Joint Data Generation. The second dataset incorporates both contextual and parametric knowledge. Here, we begin with LLaVA’s original detailed captions and annotate objects from the parametric group with special tokens. Specifically, we enclose the “omitted” objects with brackets. Formally, if denotes the original image caption sentence and represents a set of undetected objects, our data processing can be represented as:
@@
3.2 Hallucination Control
Inspired by LM-Switch (Han et al., 2023), we address hallucination by adding a control parameter , e.g., for permitted imagination and for restricting imagination, as depicted in Figure 2(a). Let represent the LLM with fixed parameters: where stands for LM-head and is the output word embedding from LM-backbone. We modify , thus making output word embedding , leading to the derived model as:
The learned projector can be regarded as the transformation from a generic word space to the object sensitive word space, where word-word semantic correspondence is optimized to object correspondence, and governs the intensity of imagining related objects.
@@ 3.2 Hallucination Control
Inference. At the inference stage, can adopt any value within the interval . Specifically, an value of corresponds to minimal reliance on parametric knowledge, whereas a value of indicates a strong inclination towards such knowledge. Detailed theoretical explanation on why HallE-Control works are elaborated upon in the Appendix.
@@ 4 Experiment
| Setting | LLM | Resolution | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object |
@@
| Control value | LLM | Resolution | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object |
@@ 4.1 Finetune on Parametric Joint Data
Before we present experiments on HallE-Control, we show the upper-bound experiments on how well the model can indicate parametric knowledge. We directly finetune the LLaVA model on parametric joint data. Intuitively, the model is trained on data indicating parametric knowledge. Its output should identify parametric knowledge accurately. Specifically, the model should put a bracket around every ”guessed” objects for indication of hallucination.
Therefore, we evaluate the object hallucination in three different settings: 1. Evaluation only on indicated objects: We do CCEval only on objects inside the bracket. The result should reflect a high level of hallucination. 2. Evaluation without indicated objects: We disregard objects in bracket and calculate CCEval. The result should reflect a low level of hallucination. 3. Evaluation with indicated objects: We calculate CCEval on all objects. Due to modification of the settings, we slightly change definition of CHAIR scores in CCEval as detailed in Appendix.
Evaluation only on indicated objects. The hallucination level for indicated objects, , is 63.90 for LLaVA7B and 62.31 for LLaVA13B. It is considerably higher than the baselines all other models. Concurrently, their coverage is 12.01 and 19.09 for LLaVA7B and LLaVA13B, respectively, which both of them significantly lower than the coverage of 33.58 for LLaVA7B baseline model. The experiment results show the object within special tokens has significantly higher hallucination rate and lower coverage rate which support our assumption that the object inside the indication tokens are objects from parametric knowledge.
Evaluation without indicated objects. For objects outside of the special token scope, we find hallucination is markedly reduced. CHAIRs decreased from 82 to 57 which is 30.5% percent improvement; CHAIRi decrease 32.4%, from 25.3 to 17.10, compared to the baseline. This suggests that the model is less prone to make erroneous assumptions for objects not marked by brackets. This is interesting because the model perfectly captures the intention of marking parametric objects in training data and replicate the behavior during inference.
Evaluation with indicated objects.: We observe a significant decline in the hallucination without any reduce in object coverage. LLaVA7B CHAIRi improved from 25.30 to 14.00 which has 44.66% improvements. For LLaVA13B CHAIRi improved from 16 to 9.86 which also has 38.38% improvements.
@@ 4.2 Hallucination Controlling
During model inference, we select 4 different , ranging from to . As shown in Table 8, we evaluate HallE-Control7B and HallE-Control13B model, which use LLaVA as backbone. For the 7B model, we train it by removing the special token in parametric joint data, showing that indication is not a necessary condition for control module to work. means the control module trained purely on contextual only data, which try to minimize the hallucination, where maximize parametric knowledge output. The results show that as increase, CHAIRi increase from 20.90 to 26.6, the coverage keeps at a similar level.
For the 13B model, we keep the indication inside the parametric joint data. The HallE-Control13B achieves the best in object existence hallucination metric. With control value set to -1 and indication, we have versus baseline’s 64 and vs baseline’s 16, and the coverage of the model is not decreasing.
@@ 5 Related Work
Large Multimodal Models (LMMs).
The rapid advancements in Large Language Models (LLMs) (Touvron et al., 2023a; Chung et al., 2022; Touvron et al., 2023b; Anil et al., 2023; Driess et al., 2023; Scao et al., 2022; OpenAI, 2023) combined with a surge in open-source initiatives, have paved the way for the emergence of extensive vision-language models (Liu et al., 2023d; Goyal et al., 2017; Zhu et al., 2023; Sun et al., 2023; Ye et al., 2023; Bai et al., 2023; Chen et al., 2023; Peng et al., 2023). LLaVA introduces the concept of integrating a simple projector during LLM fine-tuning. Chatspot (Zhao et al., 2023b) follow LLaVA’s model structure, but embed region of interest into instruction data. GPT4RoI (Zhang et al., 2023a) and Shikra (Chen et al., 2023) add grounding tasks to LLaVA structure models. Instead of using detector to provide region information to the model, we use detector to filter objects for alignment between vision and language information. Additionally, BLIP2 (Li et al., 2023d) and InstructBLIP (Dai et al., 2023) present Q-former-based LMMs. Multimodal-GPT (Gong et al., 2023) and Otter (Li et al., 2023b) aims to improve OpenFlamingo’s (Alayrac et al., ; Awadalla et al., 2023) directive adherence. mPLUG-Owl (Ye et al., 2023) suggests a two-step method: first train vision models, then refining the language model using techniques like LoRA. Our work utilize a linear layer to control object existence hallucination within LMMs.
Evaluation on LMMs.
The evaluation of large vision-and-language models (LMMs) (Yu et al., 2023; Liu et al., 2023b; c) is challenging due to the intricate nature of generation tasks they undertake. Some of the VQA-based benchmarks (Antol et al., 2015; Hudson & Manning, 2019; Gurari et al., 2018) require models to identify objects, colors, or quantities, while others (Liu et al., 2023e; Li et al., 2023c; Lu et al., 2022) offer multiple-choice questions. POPE (Li et al., 2023e) and MME (Fu et al., 2023) evaluate object hallucination using paired yes/no questions on existence, color, counting, OCR, and etc. While VQA-based benchmarks are cheap and straightforward, we find them cannot accurately reflect object hallucination for detailed captions. Besides VQA benchmarks, ROUGE (Lin, 2004; Elliott & Keller, 2014) use n-gram to evaluate similarity between ground truth and model inferences. CIDr (Vedantam et al., 2015) is a triplet-based method of collecting human annotations to measure consensus. CHAIR (Rohrbach et al., 2018) evaluates caption hallucination with object concepts. These methods are constraint by ground truth length or word variance and cannot clearly reflect hallucination with object coverage information. Wang et al. (2023) uses a language model to predict whether the caption exist hallucination, which is cheaper than GPT-4. Our work introduces CCEval, including CHAIR metrics, object coverage, average sentence length and number of objects to overcome limitations of previous evaluations.
@@
Hallucinations (Ji et al., 2023a; Shi et al., 2023; Lin et al., 2021) have been widely studied in traditional NLG (Ji et al., 2023b) tasks, including machine translation (Zhou et al., 2020; Lee et al., 2019), data-to-text (Rebuffel et al., 2021; Kasner & Dušek, 2022; Lee et al., 2022), summarization (Cao et al., 2022), dialogue (Dziri et al., 2022) and QA (Shuster et al., 2021). For LMMs, previous studies have been mainly focusing on object hallucination (Marino et al., 2019; MacLeod et al., 2017; Li et al., 2023a; e). POPE (Li et al., 2023e) reveals object existence hallucination may related with label distributions, such as object co-occurance. Earlier than POPE, Biten et al. (2022) balance object co-occurance to decrease hallucination. LRV (Liu et al., 2023a) finds the cause of hallucination in VQA benchmarks, especially unbalanced answer distribution and lack of negation information. Our work raise another important cause: misalignment between the vision and language information captured by models. More interestingly, we can explain balancing labels in Biten et al. (2022) as trying to weaken the parametric knowledge caused by the misalignment.
@@ 6 Conclusion
In summary, this study delves deep into the object hallucination phenomena within the detailed captions of LMMs, advancing understanding of the accuracy and unwarranted inference in describing visual details. We introduce a novel and comprehensive evaluation method for object existence hallucination in detailed captions. We conduct an in-depth and component-wise analysis of LMMs, meticulously examining each element that might result in hallucination. We further identify an alignment issue between the vision encoder and the instruction data. To alleviate such hallucination, we introduce controlling parameters over LMMs to condition the inference of objects.
@@ Appendix A Appendix
A.1 Why not all Hallucination Bad?
In our study, we define ”object existence hallucination” to be a phenomenon where a description makes reference to objects that is not present in the image. However, such hallucinations, when properly harnessed, can be regarded as instances of imagination. Human beings frequently use imagination to successfully accomplish tasks, often without even realizing it. Here, we present several scenarios in Table 9, Table 10, and Table 11 in which imagination and judicious inference prove to be beneficial for various downstream applications, including robotic manipulation and content moderation, among others. Our work successfully exercise control over the extent of generalization rather than attempting to eliminate all instances of imagined objects. More importantly, our model can indicate imagined objects with [object] marker.
A.2 More Details on Analyzed Models
In Section 2, we conduct analysis on different LMMs. Here, we provide more details of each model:
A.3 Training Data Quality
@@
| Image | Caption | GT Labels |
@@
The image depicts a cozy entertainment room featuring a flat screen TV situated against the wall. A dining table can be seen next to the TV, with some bottles and a cup placed on top of it. Multiple video game controllers, including Wii remotes, are also present on the dining table, suggesting that people are using the TV to play video games. To the far right of the scene, another remote is located close to the wall. A person appears to be sitting or standing behind the dining table, likely either watching the game or waiting for their turn to play. The assortment of bottles on the table suggests that the guests are enjoying drinks during their video game session. | tv: 1 cup: 1 remote: 3 bottle: 4 person: 1 dining table: 1 |
@@
In the image, an orange mass transit trolley is making its way through a city. A person is crossing the street in front of the trolley while holding a garbage bag, appearing to be cautious about the approaching vehicle. Another person is standing close to the person crossing the street, and there are two more individuals nearby. In the scene, various vehicles surround the trolley, including cars, buses, and trucks. One of the cars is parked right behind the trolley, while another is situated farther back. Two buses can be seen, with one staying behind the trolley and the other on the right side of it. A truck is also present at the far left side of the scene. A woman nearby is holding a handbag, completing the busy urban setting. | car: 3 bus: 2 truck: 1 train: 1 person: 5 handbag: 1 |
@@
The image shows a table in a restaurant, which appears to have been recently used for a meal. The table is set with four place settings, including white dishes. On the table, there are multiple cups, a bowl, and a mix of silverware like forks and spoons laid out. Some of the dishes look dirty, with used napkins and eating utensils scattered around. In the background, two people can be seen, one situated on the left side and another on the right side behind the table. There are also chairs located near the table, with one chair positioned close to the left side and another chair closer to the right side of the frame. | cup: 5 spoon: 3 fork: 2 person: 2 chair: 2 bowl: 1 dining table: 1 |
A.4 CCEval
In this paper, we introduce CCEval, a GPT-4 assisted Caption Concept Matching Evaluation. CCEval provides concept matching method and metric that can be extended to larger scale. Here, we detail the process of evaluation:
Step 0. Collect ground truth objects from annotations in a list.
Step 1. Run the model through the evaluation set of 100 images from Visual Genome. Specifically, prompt the model with ”Describe the image in detail” and obtain 100 captions as model outputs.
Step 2. Extract objects mentioned in 100 captions using GPT-4.
Step 3. For each caption, match the extracted objects with ground truth objects using GPT-4.
GPT-4 prompts are provided in the next section. After Step 3, we get the following information: matched objects (no hallucination), unmatched objects (hallucination), total objects mentioned, total ground truth objects. Our metric calculates:
Note that we do not strictly enforce consistent constraints (same coverage, average length, and average objects) across models, neither this is realistic to achieve. Although our intention is to make the model comparable and push the model to express more objects, it is sometimes not possible for some of the models, such as BLIP2, to express detailed captions, and our coverage, average length, and average number of object metrics in CCEval can accurately reflect if any model has such limitations.
In the main text, ”consistent constraints” are mainly achieved through prompting. We use the same prompt - “describe this image in detail” - across all four models in Table 2. Simply prompting can result in similar length and number of objects captions in this case since the instruction finetuning data for detailed captioning all comes from a similar source (LLaVA synthetic data). For models with different instruction finetuning data, we suggest further adjustment of the prompt or set the min_length and max_length in the inference function.
A.5 Prompt
@@ A.7 Experiment Settings
Hallucination control only involves only finetuning a linear layer added before layer. We freeze all other layers and only finetune control layer with 3 epochs. The dataset is generated data with associated value. We have 10K contextual only detailed caption data and 23K parametric joint detailed caption data.
A.8 More Experiments on LLaVA 1.5
| Setting | LLM | Resolution | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object |
| 665K baseline | LLaVA | 336x | 66.00 | 17.84 | 34.43 | 103.63 | 8.35 |
| 665K baseline | LLaVA | 336x (SW) | 64.00 | 15.54 | 34.39 | 101.33 | 8.30 |
| w/o ind. | LLaVA | 336x | 54.00 | 11.24 | 34.00 | 103.62 | 6.76 |
| w/ ind. | LLaVA | 336x | 54.00 | 8.91 | 35.94 | 103.62 | 8.53 |
The experiments on the main text are based on LLaVA 1.3, we provide experiments on LLaVA 1.5 in Table 16. The results are consistent with the conclusion in the main text and hallucination is less in this new version.
A.9 Modified CCEval of Section 4.1
1. Evaluation only on indicated objects
2. Evaluation without indicated objects
3. Evaluation with indicated objects
A.10 Qualitative Results of HallE-Control
This section shows qualitative results for HallE-Control in Table 17, Table 18, Table 19, Table 20, and Table 21. means no imagination and means max imagination. Both models can output indication.
A.11 Theoretical Explanations
HallE-Control Formulation
HallE-Control follows LLaVA’s training strategy, which freezes vision encoder and language model, only finetune the projector for the first stage. The visual encoder transfer image to visual features: